It takes 4 minutes to boil a quantity of water electrically, How long will it take to boil same quantity of water using the same heating coil but with the current doubled. Neglect any external heat losses
1
Expert's answer
2020-12-06T17:27:45-0500
"Q=Pt=P't'=I^2Rt=I'^2Rt'\\t'=\frac{I^2}{I'^2}t=\frac{1^2}{2^2}4=1\ min"